# Pervasive impact of spatial dependence on predictability

### v1

@@ 2.1 Prediction performance and spatial effect
The concept of pan-spatial statistic is implemented in analyzing the relationship between spatial dependence and spatial prediction accuracy for six datasets, comprising three sets of environment variables and three sets of socio-economic variables. The environmental data consist of distributions of three trace elements (Cu, Pb, and Zn) in a region of Australia. The socio-economic data represent the distribution of homelessness rates in three Australian cities: Melbourne, Sydney, and Brisbane. Detailed data distributions can be found in Supplementary Figure A1.
@@ 2.2.2 Performance in Non-geospatial models
In Appendix Figure 3, a scatter plot of the true values versus the predicted values for each observation in leave-one-out cross-validation is displayed. The OK method demonstrates the poorest performance in geospatial contexts, with $R^{2}$ values ranging from 0.03 to 0.26. Its predictive range is confined to a narrow interval, not approximating the actual spatial distribution. This indicates that the spatial data in these six cases are insufficient to support the construction of a reasonable semivariogram. However, by transforming the geographic space into an optimal attribute space, there is a significant improvement in predictive performance. For environment variables, the enhancements in predictive performance for the three heavy metals Cu, Pb, Zn are 0.17, 0.12, 0.24, respectively. Regarding economic variables, improvements on the homeless rate in Melbourne, Sydney, Brisbane are 0.30, 0.36, 0.09, respectively. RK, as a method that employs all explanatory variables for trend surface fitting, shows a significant improvement over OK in geographic space. Yet, in most cases, its performance is comparable to that of OK using only the two optimal spatial dimensions. This underscores the importance of spatial dependence for the predictive performance of geospatial models. Moreover, in the optimal space, RK’s predictions for the homeless rate in Sydney and Brisbane even outperform RF. In summary, the scatter plot results demonstrate that transforming geographic space into a attribute space with stronger spatial dependence allows geospatial models to more accurately capture the characteristics of spatial data, thereby making the model’s output range closer to the true data distribution and improving predictive performance.
@@ 4.1 Research data
In this study, we have meticulously gathered and analyzed six key datasets to investigate the impact of spatial dependence on spatial predictions (Appendix Figure 1). Specifically, our datasets encompass detailed distributions of trace elements within a region in Australia, as well as the distribution of homelessness rates across three major cities—Brisbane, Melbourne, and Sydney.
For the distribution of trace elements, we focused on the spatial distribution of Cu, Zn, and Pb. These elements are widely considered significant indicators of environmental pollution and are crucial for assessing environmental health. To explain the spatial variability of these trace elements, we collected nine environment-related explanatory variables, including distance to lithology (Dlith), distance to fault (Dfault), slope, water distribution, Normalized Difference Vegetation Index (NDVI), distance to main roads (MainRd), road network density (Road), soil organic carbon (SOC), and soil pH value. These explanatory variables are deemed essential for understanding and predicting the distribution of trace elements. Detailed distributions of these data can be found in the Supplementary Material Appendix Figure 1.
@@
Appendix Figure 2 displays the statistical distribution and Q-Q plots of six sets of data after logarithmic transformation. These six sets of data exhibit skewness issues in their original states, indicating deviations from a normal distribution. However, once subjected to logarithmic transformation, the distributions of these data tend to become more normal. It is important to note that, compared to the trace element data, the data on the risk of homelessness in families seem to have more outliers. Particularly in Brisbane, we can clearly observe the presence of an outlier, which may be attributed to the socioeconomic attribute data in that region being influenced by more anthropogenic factors.

### v2

@@ 2.1 Prediction performance and spatial effect
The concept of pan-spatial statistic is implemented in analyzing the relationship between spatial dependence and spatial prediction accuracy for six datasets, comprising three sets of environment variables and three sets of socio-economic variables. The environmental data consist of distributions of three trace elements (Cu, Pb, and Zn) in a region of Australia. The socio-economic data represent the distribution of homelessness rates in three Australian cities: Melbourne, Sydney, and Brisbane. Detailed data distributions can be found in Supplementary Figure S1.
@@ 2.2.2 Performance in Non-geospatial models
In Supplementary Figure S3, a scatter plot of the true values versus the predicted values for each observation in leave-one-out cross-validation is displayed. The OK method demonstrates the poorest performance in geospatial contexts, with $R^{2}$ values ranging from 0.03 to 0.26. Its predictive range is confined to a narrow interval, not approximating the actual spatial distribution. This indicates that the spatial data in these six cases are insufficient to support the construction of a reasonable semivariogram. However, by transforming the geographic space into an optimal attribute space, there is a significant improvement in predictive performance. For environment variables, the enhancements in predictive performance for the three heavy metals Cu, Pb, Zn are 0.17, 0.12, 0.24, respectively. Regarding economic variables, improvements on the homeless rate in Melbourne, Sydney, Brisbane are 0.30, 0.36, 0.09, respectively. RK, as a method that employs all explanatory variables for trend surface fitting, shows a significant improvement over OK in geographic space. Yet, in most cases, its performance is comparable to that of OK using only the two optimal spatial dimensions. This underscores the importance of spatial dependence for the predictive performance of geospatial models. Moreover, in the optimal space, RK’s predictions for the homeless rate in Sydney and Brisbane even outperform RF. In summary, the scatter plot results demonstrate that transforming geographic space into a attribute space with stronger spatial dependence allows geospatial models to more accurately capture the characteristics of spatial data, thereby making the model’s output range closer to the true data distribution and improving predictive performance.
@@ 3 Discussion and Conclusion
In spatial prediction tasks, mapping data to a feature space with strong spatial autocorrelation can facilitate the simplification of decision boundaries in prediction models, thereby enhancing prediction and classification efficiency. As illustrated in the upper part of Supplementary Figure S4, a positively autocorrelated feature space simplifies decision boundaries, which enhances the accuracy and stability of models on such data. Conversely, as demonstrated in the lower part of Supplementary Figure S4, a negatively autocorrelated feature space with complex and variable decision boundaries may increase the difficulty for models to process such data. The strength and direction (positive or negative) of spatial autocorrelation significantly influence the relationships among data points, thus determining the complexity of model learning.
In practical applications, a deep understanding and effective utilization of these spatial characteristics are crucial for designing prediction models that are tailored to specific data features. Our research reveals that the impact of geographic and non-geographic spaces on the performance of spatial predictions is universal. However, this does not imply that the characteristics of geographic data can be overlooked. Indeed, our analysis (see Figure 1) confirms that environmental variables, in particular, exhibit the strongest spatial dependencies in geographic spaces, indicating that geographic features can significantly assist prediction models in delineating clear decision boundaries.
@@ 4.1 Research data
In this study, we have meticulously gathered and analyzed six key datasets to investigate the impact of spatial dependence on spatial predictions (Supplementary Figure S1). Specifically, our datasets encompass detailed distributions of trace elements within a region in Australia, as well as the distribution of homelessness rates across three major cities—Brisbane, Melbourne, and Sydney.
For the distribution of trace elements, we focused on the spatial distribution of Cu, Zn, and Pb. These elements are widely considered significant indicators of environmental pollution and are crucial for assessing environmental health. To explain the spatial variability of these trace elements, we collected nine environment-related explanatory variables, including distance to lithology (Dlith), distance to fault (Dfault), slope, water distribution, Normalized Difference Vegetation Index (NDVI), distance to main roads (MainRd), road network density (Road), soil organic carbon (SOC), and soil pH value. These explanatory variables are deemed essential for understanding and predicting the distribution of trace elements. Detailed distributions of these data can be found in the Supplementary Figure S1.
@@
Supplementary Figure S2 displays the statistical distribution and Q-Q plots of six sets of data after logarithmic transformation. These six sets of data exhibit skewness issues in their original states, indicating deviations from a normal distribution. However, once subjected to logarithmic transformation, the distributions of these data tend to become more normal. It is important to note that, compared to the trace element data, the data on the risk of homelessness in families seem to have more outliers. Particularly in Brisbane, we can clearly observe the presence of an outlier, which may be attributed to the socioeconomic attribute data in that region being influenced by more anthropogenic factors.
@@ 4.2 Experiment design
This study treats spatial prediction as a general term that encompasses both geographic interpolation and extrapolation. We investigate how generalized spatial prediction methods perform in relation to both environmental and socio-economic variables. To achieve this, we selected trace elements as an example of environmental variables and homelessness as an example of socio-economic variables. We intentionally chose not to differentiate between datasets based on whether they are discrete or continuous, focusing instead on their spatial characteristics, allowing us to explore the broader applicability of spatial prediction models across different fields and data types.
@@
It should be mentioned that Moran’s I index are normally applied to discrete spatial data. For both trace elements and homelessness data, we consider the datasets to be spatially discrete. As such, using the Moran’s I index to assess their spatial autocorrelation is appropriate. Specifically, for predicting trace elements, applying Moran’s I to analyze their spatial patterns is a common and valuable practice. For instance, LISA can help identify hotspots in the spatial distribution of a specific metal concentration. While geostatistics models are frequently applied to trace elements, it is uncommon to see them used for predicting socio-economic attributes such as homelessness. However, in many real-world applications, census-derived statistical data is often used to fill in missing values, especially when there are insufficient explanatory variables for a regression-based prediction model. In these cases, interpolation techniques, including geostatistics, are often employed.